OpenStudy (anonymous):

Hi i'm trying to verify that 0v=0 is a linear map where the zero on the left hand side is the the function

OpenStudy (anonymous):

So your function is f(v)= 0*v. For any v and w: \[f(v + w) = 0*(v+w) = 0 = 0*v + 0*w = f(v) + f(w)\] That proves the first condition of a linear map. You should be able to do the second.

OpenStudy (anonymous):

idk.